OpenStudy (cutiecomittee123):

Which of the following are identities Check all that apply sin^2x-cos^x=1 sin^2x=1-cos^2x tan^2x=sec^2x-1 cot^2x+csc^2x=1

OpenStudy (jacob902):

the correct identities are: 1 + tan^2x = sec^2x 1 + cot^2x = csc^2x sin^2x + cos^2x = 1

OpenStudy (jacob902):

@cutiecomittee123 you blocked me ?

OpenStudy (anonymous):

\[\sin ^2x+\cos ^2x=1\] \[\sec ^2x-\tan ^2x=1\] \[\csc ^2x-\cot ^2x=1\] rearrange and see which apply.

OpenStudy (cutiecomittee123):

All three of those work because sin^2x=1-cos^2x tan^2x=sec^2x-1 cot^2x+csc^2x=1

OpenStudy (cutiecomittee123):

wait only the first two. not the last one that I listed

OpenStudy (anonymous):

correct.